### v1

# Efficient 3D Bayesian Full Waveform Inversion and Analysis of Prior Hypotheses

Abstract
Spatially 3-dimensional seismic full waveform inversion (3D FWI) is a highly nonlinear and computationally demanding inverse problem that constructs 3D subsurface seismic velocity structures using seismic waveform data. To characterise non-uniqueness in the solutions we demonstrate Bayesian 3D FWI using an efficient method called physically structured variational inference applied to 3D acoustic Bayesian FWI. The results provide reasonable posterior uncertainty estimates, at a computational cost that is only an order of magnitude greater than that of standard, deterministic FWI. Furthermore, we deploy variational prior replacement to calculate Bayesian solutions corresponding to different classes of prior information at low additional cost, and analyse those prior hypotheses by constructing Bayesian L-curves. This reveals the sensitivity of the inversion process to different prior assumptions. Thus we show that fully probabilistic 3D FWI can be performed at a cost that may be practical in small FWI problems, and can be used to test different prior hypotheses.
1 Introduction
Seismic full waveform inversion (FWI) uses both phase and amplitude information from observed waveform data to estimate compatible subsurface seismic velocity maps [7]. Since FWI is a high dimensional, strongly nonlinear and ill-posed inverse problem, finding the true Earth model and its corresponding uncertainties using standard inversion methods is currently impossible [2, 17, 10].
Monte Carlo sampling methods are commonly used to solve nonlinear Bayesian inverse problems and quantify uncertainties. However, applying conventional Monte Carlo methods to Bayesian FWI is computationally expensive due to the curse of dimensionality [3]. Variational inference uses optimisation rather than random sampling to solve Bayesian problems, and can be more efficient and easier to scale to high dimensions [16]. Recently, variational methods have been applied to reduce the computational cost of both Bayesian 2D [9, 13] and 3D FWI [11, 6], to reduce the computational cost.
We demonstrate Bayesian 3D FWI with substantially improved efficiency compared to previous methods by using physically structured variational inference [PSVI: 14]. This method has proven to provide relatively accurate inversion results including uncertainty estimates at greatly reduced computational cost in 2D FWI problems. To analyse, test and potentially select between different prior hypotheses, we use variational prior replacement [VPR: 12] to obtain inversion results corresponding to different prior pdf’s at negligible additional computational cost.
In what follows we first summarize methods, then present 3D FWI results. We compare inversion results obtained using different prior hypotheses by building Bayesian L-curves, then discuss implications of this work and draw conclusions.
2 Methodology
2.1 Variational Bayesian Inversion
Bayesian inference calculates the posterior probability distribution function (pdf) of model parameters $\mathbf{m}$ given observed data $\mathbf{d}_{obs}$ using Bayes’ rule
$p(\mathbf{m}|\mathbf{d}_{obs})=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p(\mathbf{m})}{p(\mathbf{d}_{obs})}$ (1)
where $p(\mathbf{m})$ represents prior information about $\mathbf{m}$, and $p(\mathbf{d}_{obs}|\mathbf{m})$ is the likelihood of observing $\mathbf{d}_{obs}$ given any value of $\mathbf{m}$. Term $p(\mathbf{d}_{obs})$ is a normalisation constant called the evidence.
In variational inference, an optimal variational distribution $q(\mathbf{m})$ is selected from a family of distributions to best approximate the posterior distribution $p(\mathbf{m}|\mathbf{d}_{obs})$. This is often accomplished by minimising the Kullback-Leibler (KL) divergence (a measure of difference) between the posterior and variational distributions [5], or equivalently maximising the evidence lower bound (ELBO) of $\log p(\mathbf{d}_{obs})$ defined as:
$\text{ELBO}[q(\mathbf{m})]=\mathbb{E}_{q(\mathbf{m})}[\log p(\mathbf{m},\mathbf{d}_{obs})-\log q(\mathbf{m})]$ (2)
This is an optimisation problem with a fully probabilistic result.
2.2 Physically Structured Variational Inference (PSVI)
PSVI is an efficient variational method that employs a transformed Gaussian distribution with a specific covariance structure to best fit the posterior distribution [14]. A Gaussian distribution $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ is characterised by a mean vector $\boldsymbol{\mu}$, and a covariance matrix $\boldsymbol{\Sigma}$ which is often re-parametrised using Cholesky factorisation $\boldsymbol{\Sigma}=\mathbf{L}\mathbf{L^{T}}$, where $\mathbf{L}$ is a lower triangular matrix.
PSVI captures only specific (important) posterior correlations in parameters $\mathbf{m}$, guided by prior knowledge of the physics controlling typical similar inverse problems. Specifically, our implementation of PSVI includes posterior correlations between pairs of locations that are spatially proximal, typically within one dominant wavelength in FWI problems. All other parameter correlations are ignored. This is accomplished by solving for specific off-diagonal elements of $\mathbf{L}$ during inversion, and fixing all other elements to be zero. This sparse structure for $\mathbf{L}$ and hence $\boldsymbol{\Sigma}$, requires a manageable number of hyperparameters to be optimised while still capturing what are typically the most significant correlations.
Considering the computational cost of 3D Bayesian FWI, a top priority is to improve its efficiency. PSVI does so at the cost of some loss of generality in the solution found due to the Gaussian foundations for the posterior distribution. [14] demonstrated that the method nevertheless produces reasonable statistical information about the full, nonlinear posterior distribution.
2.3 Variational Prior Replacement
Say we have two different prior pdf’s $p_{old}(\mathbf{m})$ and $p_{new}(\mathbf{m})$ (hereafter subscripts old and new are used to denote the order in which the pdf’s are used). According to Bayes’ rule (equation 1), the two posterior probability distributions can be calculated by:
$p_{old}(\mathbf{m}|\mathbf{d}_{obs})=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p_{old}(\mathbf{m})}{p_{old}(\mathbf{d}_{obs})}\approx q_{old}(\mathbf{m})$ (3)
and
$\begin{split}p_{new}(\mathbf{m}|\mathbf{d}_{obs})&=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p_{new}(\mathbf{m})}{p_{new}(\mathbf{d}_{obs})}\\ &=\dfrac{p(\mathbf{d}_{obs}|\mathbf{m})p_{old}(\mathbf{m})}{p_{old}(\mathbf{d}_{obs})}\ \dfrac{p_{new}(\mathbf{m})}{p_{old}(\mathbf{m})}\ \dfrac{p_{old}(\mathbf{d}_{obs})}{p_{new}(\mathbf{d}_{obs})}\\ &=k\ p_{old}(\mathbf{m}|\mathbf{d}_{obs})\ \dfrac{p_{new}(\mathbf{m})}{p_{old}(\mathbf{m})}\\ &\approx k\ q_{old}(\mathbf{m})\ \dfrac{p_{new}(\mathbf{m})}{p_{old}(\mathbf{m})}\\ &\approx q_{new}(\mathbf{m})\end{split}$ (4)
In equation 3, we assume that the Bayesian problem given $\mathbf{d}_{obs}$ and prior pdf $p_{old}(\mathbf{m})$ is solved using variational inference giving an approximate posterior $q_{old}(\mathbf{m})$. The third and forth lines in equation 4 are obtained by substituting equation 3 into the second line. Term $p_{new}(\mathbf{m})/p_{old}(\mathbf{m})$ updates the solution that includes the old prior to a solution that contains the new one, and we set constant $k=p_{old}(\mathbf{d}_{obs})/p_{new}(\mathbf{d}_{obs})$. This implies that the new posterior distribution can be calculated from the old one by updating prior information post inversion by a simple multiplication [8].
Evaluating $k$ requires the two evidence terms to be calculated which is computationally intractable. Zhao and Curtis [12] estimated a second variational distribution $q_{new}(\mathbf{m})$ to approximate $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$ (fifth line in equation 4) by minimising the KL divergence between the new posterior and $q_{new}(\mathbf{m})$, without requiring the value of $k$ to be known explicitly. This step is called variational prior replacement (VPR), and we use PSVI to perform VPR because its probability value is cheap and easy to evaluate for any model parameter values post inversion. Using PSVI to find $q_{old}(\mathbf{m})$ thus allows us to perform VPR efficiently to find $q_{new}(\mathbf{m})$ in equation 4 for alternative prior distributions.
3 3D Bayesian FWI Example
We apply PSVI to a synthetic 3D acoustic FWI problem. The true velocity model is a part of the 3D overthrust model, which contains $63\times 101\times 101$ grid cells in Z (depth), Y and X directions, respectively, with a cell size of 50 m in each direction (Figure 1a). We model 81 impulsive surface sources (red stars) and 10,201 receivers (white dots) with spacings of 500 m and 50 m, respectively: details in Appendix A.
We define the same Uniform prior distribution as Zhang et al. [11], and define an uncorrelated diagonal Gaussian likelihood function for variational Bayesian inversion (details in Appendix A.). The inversion is performed using 6 s long waveform data in bands referred to as the low (1 – 6 Hz), intermediate (1 – 10 Hz), and high frequency band (1 – 13 Hz). Optimizations for posterior pdf’s using data in intermediate and high frequency bands used results from the previous band as a starting point.
3.1 Inversion Results
Figures 1a - 1d display the inversion results obtained using the high frequency band data; results using low and intermediate frequency bands are provided in Appendix A. From left to right, each figure shows the true velocities, the posterior average velocity, standard deviation, and relative error maps, where the relative error is the absolute difference between true and mean velocities divided by the standard deviation at each point. The top two rows show vertical sections Y = 2.5 km and X = 2.5 km; bottom row shows a horizontal section Z = 1.25 km.
The three mean sections resemble the true velocity structures, and the overall relative errors are less than 3, indicating that differences between the true and inverted mean models are within three standard deviations, as would be expected of the true solution to this Bayesian problem.
Figure 1: (a) - (d) High frequency 3D variational Bayesian FWI results obtained using PSVI and the Uniform prior distribution. (e) - (h) Inversion results after fine tuning the VPR results, as explained in the Discussion. In each figure, top two rows show vertical sections Y = 2.5 km and X = 2.5 km; bottom row shows a horizontal section Z = 1.25 km.
3.2 Testing Different Prior Hypotheses
In addition to the Uniform prior distribution, we introduce a set of smoothed prior distributions with different strengths of spatial smoothness controlled by a hyperparameter $\sigma_{s}$ (larger $\sigma_{s}$ allows rougher models - see Zhao and Curtis [12]). We consider 7 smoothed prior pdf’s with different $\sigma_{s}$ values: 2000, 1000, 500, 250, 125, 62.5 and 31.25, respectively. In the later part of this paper, we also denote the Uniform prior distribution as $\sigma_{s}=\infty$ (no smoothing). These 8 prior pdf’s are used to mimic a situation in which we have different prior hypotheses and wish to analyse their consequences.
The posterior pdf’s using each of the 7 smoothed priors are obtained by replacing the Uniform prior pdf in inversion results from the previous section (Figures 1a to 1d) by the respective smoothed prior pdf’s, using VPR. Thus, no additional inversion is performed to obtain these results, and in Appendix B, we verify their quality by comparing them to results obtained by performing independent Bayesian inversions. Figure 2 displays the results using all 8 prior pdf’s on vertical section Y = 2.5 km, each row presenting the posterior pdf obtained using the $\sigma_{s}$ value indicated on the left. The posterior marginal pdf’s of two vertical profiles at locations X = 1 km and X = 3 km are displayed in Appendix C.
Figure 2: Inversion results obtained using (a) the Uniform prior pdf, and (b) – (h) the 7 smoothed priors with decreasing values of $\sigma_{s}$ (increasing strength of smoothness), in vertical section Y = 2.5 km.
As the magnitude of smoothness increases from Figures 2a to 2h, both the posterior samples and the mean velocity maps exhibit fewer spatial variations. The standard deviations decrease (note that colorbar scales vary in the third column); this leads to increased relative errors wherever (true) large velocity contrasts between neighbouring cells are precluded by the smoothing prior information for small $\sigma_{s}$.
By comparing each set of results with the true velocity model in Figure 1a, we find that smoothed prior distributions with $\sigma_{s}=1000$ or $\sigma_{s}=500$ (Figures 2c and 2d) are relatively good choices for this FWI problem. However, in reality we never know the true velocity structure, so can not calculate or use relative errors to choose between prior distributions. In deterministic FWI, an order of preference is often constructed by calculating data misfit values arising from the different inversion results (referred to as an L-curve). Following a similar approach, we compare the 8 posterior distributions by drawing N = 100 samples from each posterior pdf and performing forward simulations to calculate corresponding misfit values between the observed and synthetic data. This results in 8 approximate probability distributions over misfit values from the inversion results using different priors, displayed in Figure 3a. As expected, the overall misfit values increase as more smoothness is injected into the inversion.
The mean and standard deviation of each distribution are displayed in Figure 3b by black dots and error bars, and are connected by a red line which we call a Bayesian L-curve. From left to right along the curve, the magnitude of prior information (relative smoothness in this case) decreases from $\sigma_{s}=31.25$ (top-left point) to $\sigma_{s}=\infty$ (no smoothing – bottom-right point). An arrow marks the case with $\sigma_{s}=1000$, and we observe that data misfits start to decrease significantly if we further increase the roughness value $\sigma_{s}$ of the prior information, implying that this value provides a balance between information from the smoothing distribution and the observed data.
More generally, many different types of prior hypotheses can be represented by a hypothesis space $\mathcal{H}$ represented schematically in Figure 3c. Given a rather non-informative Uniform prior distribution in the old posterior pdf, in principle we can inject any additional prior assumptions into the inversion; each would have different effects on the inversion results, between which we may wish to discriminate. As an example, we consider another prior assumption using the real-geology informed prior distribution from Zhao and Curtis [12] - described in Appendix D. The dashed blue line in Figure 3b displays the corresponding Bayesian L-curve for this set of prior assumptions, where the strength of geological prior information decreases from left to right. The data misfits are lower than those from smoothing priors in red because geologically informed priors impose less smoothness vertically than horizontally (which is correct in the true model - and often in the Earth).
Figure 3: (a) Probability density functions for data misfit values calculated from posterior pdf’s displayed in Figure 2 using differently smoothed priors. (b) Bayesian L-curves representing average data misfits and their standard deviations (error bars) with respect to different prior hypotheses. The horizontal axis represents the normalised (relative) weakness of prior information applied by each prior distribution: the value 1 represents the Uniform prior pdf. The red line denotes the L-curve using smoothed prior pdf’s, and dashed blue line denotes the equivalent curve when adding the geological prior information. (c) A space $\mathcal{H}$ containing all possible prior hypotheses. Black dot represents a non-informative Uniform prior distribution, and we denote different prior assumptions such as smoothed and geological prior pdf’s in this study by different paths through $\mathcal{H}$.
3.3 Computational Cost
The inversion using the Uniform prior distribution (Figures 1a to 1d) is obtained by performing 5200 forward and adjoint simulations. To further reduce the computational cost, we use a minibatch of 36 shot gathers, randomly selected from the total of 81 shots in each FWI simulation [11]. Note that in deterministic FWI, one would usually perform $\sim$100 iterations to obtain a reasonable result, and a full batch of 81 shots is often used at each iteration. Since forward simulation using 81 shot gathers would require 2.25 times more computation than that using a minibatch of 36 shots, the computational cost of 100 iterations implemented in deterministic FWI is equivalent to 225 forward evaluations performed in this study.
With PSVI we therefore obtain reasonable probabilistic uncertainty estimates using only an order of magnitude more computation than deterministic FWI, from which no robust uncertainty information can be obtained. In addition, this single set of variational inversion results allows multiple posterior distributions corresponding to different prior pdf’s to be produced using VPR, without any further FWI simulations [12].
4 Discussion
Rather than approximating the true solution to the desired inverse problem directly, the VPR approximation to each new posterior pdf includes only information about the true solution that is already included within in the old posterior pdf. It therefore carries over an imprint of any errors in the latter [12]. VPR results may therefore differ from independent Bayesian inversion results. To improve the results, we can refine (fine tune) the outcomes obtained from VPR using observed data. We do this by performing another variational Bayesian inversion using PSVI, but with a smaller number of iterations which start from the VPR output, because the VPR solution should already be reasonably close to the true posterior pdf. Figures 1e to 1g show the fine-tuned inversion results with 800 additional simulations, which are improved compared to the VPR results (Figure 7 in Appendix B). To reduce this additional computation we might again depart from formal Bayesian analysis by using the Bayesian L-curve obtained from VPR to select a smaller subset of prior hypotheses to entertain, then fine tune only those VPR posterior pdf’s.
5 Conclusions
We solve a 3D Bayesian FWI problem using physically structured variational inference, in which a transformed Gaussian distribution is employed to approximate the posterior distribution of the fully nonlinear inverse problem. Using a non-informative Uniform prior distribution we obtain a posterior solution with reasonable uncertainty estimates. The inversion progresses through low, intermediate and high frequency bands, with the total computational cost only an order of magnitude higher than deterministic FWI. In addition, we explore multiple different prior distributions and calculate the corresponding posterior distributions using variational prior replacement, without the need for additional FWI simulations. These results are used to analyse different prior hypotheses through the construction of Bayesian L-curves. We thus demonstrate that accurate Bayesian FWI and analysis or even a choice of prior hypotheses is in principle now feasible in 3D.
6 ACKNOWLEDGMENTS
We thank the Edinburgh Imaging Project (EIP - https://blogs.ed.ac.uk/imaging/) sponsors (BP and TotalEnergies) for supporting this research.
Appendix A Inversion results using a Uniform Prior Distribution
As displayed in Figure 4a, the true velocity model used in the main text is a part of the 3D overthrust model [1], which contains $63\times 101\times 101$ grid cells in Z (depth), Y and X directions, respectively, with a cell size of 50 m in each direction. We deploy 81 sources (red stars) and 10,201 receivers (white dots) at the surface with spacings of 500 m and 50 m, respectively. Data are generated by solving the 3D acoustic wave equation using a time-domain pseudo-analytical method [4] with an Ormsby wavelet for 6 s. We define a Uniform prior distribution for the velocity values at different depths, with upper and lower bounds shown in Figure 4b.
Figures 5 and 6, respectively, display the inversion results obtained using low and intermediate frequency waveform data, performed using physically structured variational inference [PSVI – 14]. Similarly to the inversion results presented in Figure 1 in the main text, the top two rows display two vertical sections at locations Y = 2.5 km and X = 2.5 km, and bottom row illustrates one horizontal section at Z = 1.25 km. From left to right, each column shows three true velocity sections (marked by black dashed lines in Figure 1a), the corresponding inverted average velocity, the standard deviation, and the relative error maps of the posterior distribution, where the relative error is the absolute difference between the true and mean values divided by the standard deviation at each point. Comparing these results with those using the high frequency data we observe that as the frequency of waveform data used for inversion increases, the spatial resolution becomes higher, and velocity estimates become more accurate. The overall relative errors are within 3 standard deviations, which might be expected of the true posterior probability distribution.
Figure 4: (a) True velocity model. Source and receiver locations are indicated by red stars and white dots at the surface (the latter are so close together that they may be observed as a pale haze). Dashed black lines mark the locations of three 2D slices discussed in the text. (b) Upper and lower bounds of a Uniform prior distribution at each depth.
Figure 5: Low frequency 3D variational Bayesian FWI results obtained using PSVI and the Uniform prior distribution. Top two rows show two vertical sections at Y = 2.5km and X = 2.5km, and bottom row shows a horizontal section at Z = 1.25km. Panels (a) – (d) display the true velocity model, the mean, the standard deviation and the relative error maps of the posterior distribution.
Figure 6: Intermediate frequency 3D variational Bayesian FWI results.
Appendix B Verification of Variational Prior Replacement
In this section, we test the effectiveness and accuracy of variational prior replacement [VPR – 12] for 3D Bayesian FWI by comparing the posterior solutions obtained using VPR to those obtained from independent Bayesian inversions. For VPR, we consider the Uniform prior distribution defined in the main text as the old prior pdf, which is then removed from the old posterior distribution (displayed in Figures 1a to 1d in the main text) and replaced by a smoothed prior distribution with $\sigma_{s}=1000$ (see details in the main text). The corresponding results are displayed in Figure 7. We then perform an independent variational Bayesian inversion using the same smoothed prior distribution. Similarly to the Uniform prior case, the inversion is performed using the same three frequency bands of waveform data with the same number of (5200) forward simulations. Figure 8 illustrates the results. The main features of these two results are roughly consistent, indicating that VPR can replace the old with the new prior information in the inversion results without solving a Bayesian inverse problem repeatedly from scratch. Some small discrepancies between these two sets of results remain, especially in the deeper parts of the model. This is partly because VPR introduces a variational distribution $q_{new}(\mathbf{m})$ to approximate the new posterior pdf $p_{new}(\mathbf{m}|\mathbf{d}_{obs})$, rather than calculating the true new posterior distribution, as expressed in the fifth line in equation 4 in the main text [8, 12]. In addition, this also indicates that VPR might not be as accurate for this extremely high dimensional and complex 3D Bayesian inverse problem, compared to previous tests in 2D [12]. Nevertheless, the result in Figure 7 is obtained almost for free, compared to the huge computational cost typically involved in 3D forward simulation [15] and Bayesian inversion [11]. In the main text, we introduce a fining tune scheme to improve the inversion results displayed in Figure 7.
Figure 7: Variational prior replacement results obtained by removing the Uniform prior information from the old posterior pdf, and imposing a smoothed prior pdf with a smoothness value of $\sigma_{s}=1000$.
Figure 8: Independent variational Bayesian inversion results obtained using the smoothed prior distribution with a smoothness value of $\sigma_{s}=1000$. This is used to verify VPR results displayed in Figure 7.
Appendix C Posterior Marginal pdf’s using Smoothed Priors
Figure 9 compares posterior marginal pdf’s of two vertical velocity profiles at locations X = 1km (top row of Figure 9) and X = 3km (bottom row), marked by red dashed lines in Figure 7a. Figure 9a displays the posterior marginal pdf’s obtained from the Uniform prior distribution, and Figures 9b – 9h show those from the smoothed prior pdf’s – the smoothness values $\sigma_{s}$ are noted in the title. In each figure, the red line shows the true velocity profile and black line displays the inverted mean velocity profile. We observe that with the increase of the magnitude of smoothness injected by the prior information, the posterior marginal pdf’s become smoother with fewer spatial variations (thus lower uncertainties), similarly to the conclusions drawn in the main text.
Figure 9: Posterior marginal pdf’s at two vertical locations marked by red dashed lines in Figure 7a. (a) shows results obtained using the Uniform prior pdf, and (b) – (h) correspond to those using the smoothed prior pdf’s with different $\sigma_{s}$ values.
Appendix D Inversion results using Geological Prior information
In addition to the smoothed prior distributions mentioned in the main text, we also consider another set of prior distributions with real-geology informed inter-parameter correlation information. This is achieved by selecting a set of realistic geological images and then subsampling them using a smaller window with a predefined window size. These subimages are used to calculate a local correlation matrix to define prior correlations between pairs of model parameters. More details about this geological prior distribution can be found in [12]. In this section, we consider 5 such geological prior distributions with different local window sizes of $2\times 2\times 2$, $4\times 4\times 4$, $6\times 6\times 6$, $8\times 8\times 8$ and $10\times 10\times 10$, respectively. Again, the corresponding posterior distributions are calculated using VPR by replacing the old Uniform prior distribution by each defined geological prior distribution. Figure 10 displays statistics of the corresponding posterior pdf’s, where the sizes of the correlation windows used are denoted on the left side of Figure 10. Figure 10a represents the results obtained using the Uniform prior distribution. We observe that larger prior correlation windows generally result in better inversion results that are more similar to the true velocity model, since a larger window can define better prior correlation information, which improves the inversion results. However, a larger window also has higher memory requirements. These inversion results are then used to construct a Bayesian L-curve, as displayed by the dashed blue line in Figure 3b in the main text. From left to right along this dashed blue line, the correlation window size decreases from $10\times 10\times 10$ (top-left point) to $0\times 0\times 0$ (bottom-right point), the latter being the Uniform prior pdf with no prior correlation.
Figure 10: Inversion results on a vertical section Y = 2.5 km obtained using (a) the Uniform prior pdf and (b) – (f) the geological prior pdf’s with different sizes of local correlation windows, denoted on the left side of each row. From left to right, they represent one posterior sample, the mean velocity, the standard deviation and the relative error maps.